# A Primer on SO(3) Action Representations in Deep Reinforcement Learning

### v1

@@ Abstract
Many robotic control tasks require policies to act on orientations, yet the geometry of $\mathrm{SO}(3)$ makes this nontrivial. Because $\mathrm{SO}(3)$ admits no global, smooth, minimal parameterization, common representations such as Euler angles, quaternions, rotation matrices, and Lie algebra coordinates introduce distinct constraints and failure modes. While these trade-offs are well studied for supervised learning, their implications for actions in reinforcement learning remain unclear. We systematically evaluate $\mathrm{SO}(3)$ action representations across three standard continuous control algorithms, PPO, SAC, and TD3, under dense and sparse rewards. We compare how representations shape exploration, interact with entropy regularization, and affect training stability through empirical studies and analyze the implications of different projections for obtaining valid rotations from Euclidean network outputs. Across a suite of robotics benchmarks, we quantify the practical impact of these choices and distill simple, implementation-ready guidelines for selecting and using rotation actions. Our results highlight that representation-induced geometry strongly influences exploration and optimization and show that representing actions as tangent vectors in the local frame yields the most reliable results across algorithms.
@@ 1 Introduction
What is still missing is a general, systematic evaluation of $\mathrm{SO}(3)$ representations in deep reinforcement learning (RL). While the intuitions and results from prior studies on input/observation representations for orientations also apply to the RL setting, the most suitable $\mathrm{SO}(3)$ action representation remains unclear. Action representation requires special attention, as it shapes the exploration dynamics induced by stochastic policies and exploration noise, and has implications for action clipping to comply with actuation constraints. Prior works have proposed specific action representations for a narrow set of problems and algorithms (Alhousani et al., 2023a; b). Schuck et al. (2025) recently attempted to tackle the general question of best action and observation representations for RL, but limited their investigation to Deep Deterministic Policy Gradients (DDPG) under sparse rewards.
@@ 2.1 The SO(3) Manifold
Representation Dim. Cover Smooth Singularities Constraints
Matrix $\bm{R}$} 9 single + - ${\bm{R}}^{T}{\bm{R}}=\bm{I}$, $\det{\bm{R}}=1$
Quaternion $\bm{q}$} 4 double + - ${\bm{q}}||_{2}=1$
Euler angles $(\phi,\theta,\psi)$ 3 multi - +
Lie algebra $\mathfrak{m}$ 3 multi - +
@@ 2.4 Projections
For deterministic policies, this suffices to guarantee actions lie in $\mathrm{SO}(3)$. Stochastic policies, on the other hand, are often parameterized as multivariate Gaussians. Applying the projection to each sampled action guarantees on-manifold actions, but it also warps the action distribution and renders log probabilities intractable. Algorithms such as PPO and SAC rely on accurate log probabilities, and closed-form corrections for normalization on $\mathcal{S}(3)$ or SVD-based projections are not readily available.
In this paper, we adopt a practical compromise: we project the mean inside the network wherever possible, then sample in the ambient Euclidean space and let the environment project the sampled, off-manifold action to a valid rotation. This approach keeps training compatible with standard log-probability computations while ensuring feasibility at execution time.
@@ 3.1 Environment Setup
with the geodesic distance $d({\bm{R}}_{1},{\bm{R}}_{2})=\arccos\Big(\tfrac{\mathrm{tr}({\bm{R}}_{1}^{\top}{\bm{R}}_{2})-1}{2}\Big).$ Intuitively, equation 3 takes the shortest path towards the desired orientation ${\bm{R}}_{a}$ with a maximum step length of $\alpha_{max}$. Figure 1 visualizes the environment dynamics. Dense rewards $r_{t}^{\mathrm{dense}}=-d({\bm{R}}_{t},{\bm{R}}_{g})$ are defined as the negative angle to the goal. Sparse rewards are 0 when the angle between state and goal $d({\bm{R}}_{t},{\bm{R}}_{g})\leq\frac{\pi}{10}$ and -1 everywhere else. Termination occurs after a fixed step limit of $50$.
@@ 3.2 Performance Comparison
We benchmark PPO, SAC, and TD3 in the pure-rotation environment using four action parameterizations: rotation matrices, unit quaternions, tangent-space (rotation vectors), and Euler angles, each evaluated as global and delta actions. All other components remain fixed: network architectures, training budgets, observation spaces, and reward definitions are identical across conditions. The results are presented in table 2, with the best results highlighted in blue, and the second-best shown in bold. Results are averaged over 50 runs each. See section A.2 for the complete training curves, and section A.6 for further information on hyperparameters.
@@ 3.3 Explaining the Effects of SO(3) Action Representations
Figure 2: 3D distribution of Euler angles sampled from $\mathcal{N}(0,2)$ and squashed with $\tanh$ in the Lie algebra $\mathfrak{m}$.
@@
Hypothesis 4 Scaling actions to the range of permissible rotations boosts performance. As outlined in section 2.5, tangent vector increments in the local frame can easily be scaled. Restricting the network output to the range of possible rotations should be more efficient because the agent does not need to learn that actions with the same direction and magnitudes larger than $\alpha_{max}$ lead to the same outcome. Furthermore, it removes discontinuities at the cut locus from the action space and allows more fine-grained control in the relevant action ranges.
We can see the effects in all three algorithms. Ablation studies with unscaled tangent increments and dense rewards consistently exhibit a performance difference of around 1.5-1.5 compared to scaled tangent vectors across PPO, SAC, and TD3. The effect is smallest for PPO, where the policy initialization around zero prevents frequent sampling of unscaled actions that reach the cut locus. Agents trained on sparse rewards exhibit a similar decrease in performance. More importantly, however, sparse reward agents sometimes take longer in TD3 to converge to a near-optimal policy, and may diverge completely in SAC (see section A.1.5). SAC’s failures are caused by the entropy bonus driving actions to regions of the representation with discontinuities and singularities as outlined above.
@@ 4 Robotic Benchmarks
In real-world problems, rotation actions in $\mathrm{SO}(3)$ rarely act on purely rotational dynamics as in section 3. More often, they appear in a larger context, such as controlling a robot arm’s position, orientation, and gripper to grasp objects, or controlling a drone’s orientation and total thrust for trajectory tracking. In this section, we thus validate the relevance and transferability of our findings by applying the different action parameterizations on three robot benchmarks. In this set of studies, we focus on the most promising action combinations, i.e., global matrix and quaternion representations, and delta tangent and Euler angles.
@@
In figure 3, action representations significantly impact convergence speed on both benchmarks. Actions in the local tangent space consistently outperform other representations by converging faster and achieving higher rewards. Surprisingly, Euler angles are second before rotation matrices and quaternions. The reason is the limited range of angles required in the tasks. The drone cannot deviate too much from the upright orientation without crashing, and hence the policy remains in a region of $\mathrm{SO}(3)$ where Euler angles are still well-behaved. In contrast, absolute quaternion and matrix actions are highly random at initialization, which leads to fast crashes and limits progress (see section A.1.1).
@@ 5 Conclusion
One limitation of this paper is its restriction to state-based observations and small networks. Observations do not affect the action space; thus, our results likely still apply. However, this requires empirical evidence. We also noticed a lack of suitable benchmarks that require control over the full $\mathrm{SO}(3)$ manifold. Our extension to the HER environments could act as a starting point to build up a standard set of tasks focusing on this capability. Finally, diffusion policies have found widespread adoption for imitation learning in robotics. A similar study on suitable representation choices for diffusion may reach significantly different conclusions due to diffusion models’ multi-modality and noise processes.
@@ 5.1 Reproducibility Statement
In case of acceptance, we will provide links to the code for all environments, benchmarks, and training scripts required for reproducing our results in this section. This is not possible during the review due to anonymization concerns.
@@ A.1.1 Projections of Noise Samples
For stochastic policies used in PPO and SAC, a natural question is whether projecting sampled actions to the valid representation manifold speeds up convergence, as the critic never sees off-manifold actions. Ablation runs in figure 7 show this is untrue. Projections lead to a significant performance loss in PPO, because the probability ratios $\tfrac{\pi_{\theta}(\mathbf{a}\mid\mathbf{s})}{\pi_{\theta_{old}}(\mathbf{a}\mid\mathbf{s})}$ used in its clipped surrogate objective no longer matches the ratio of the unprojected action. Newly computed probabilities use projected actions, while the stored probabilities of the old policy are based on non-projected actions. SAC remains unaffected as action probabilities are re-computed online during policy updates, with projections applied afterwards. However, we do not observe any performance improvements.
@@
Based on our ablations, we recommend not projecting the action samples of PPO and SAC, and instead relying on the environment for the action projection. Introducing projections offers no performance gains in the best case, and has the potential to severely harm convergence.
@@ A.1.5 Scaling Tangent Vectors
Section 2.5 outlines how tangent vectors can be scaled to only encompass the range of possible rotations with a limited angle $\alpha_{max}$. In section 3.3, we explain how this helps avoid the cut locus of the tangent space at large action norms. Here, we present the training curves of our ablations in figures 9, 10 and 11. As previously stated, the final mean reward of scaled policies is slightly increased by $1.5$ to $2$ due to more fine-grained control. Much of the unscaled policies’ action space lies outside the $\alpha_{max}$ limit and thus maps to the same actions.
@@
The effect is smallest for PPO, where the initialization of the policy leads to actions close to zero. For SAC and TD3, the effect is slightly more visible in dense environments. Agents in sparse environments infrequently converge with significant delay (TD3) or fail to converge at all (SAC), causing the increase in variance across runs. The entropy bonus causes the complete failures in SAC, which drives actions towards larger magnitudes around the cut locust before learning a reasonable policy. Dense rewards prevent this by balancing the entropy bonuses with a continuous signal towards successful policies from the beginning.
@@ A.1.7 Extended Results for Entropy Regularization in PPO and SAC
A.2 Complete Experimental Results
@@ A.2.2 SAC Results
For dense rewards and global actions, the rotation matrix representation converges fastest. We attribute this to the uniqueness and smoothness of the representation. Quaternions are second with decreased convergence stability because of the multi-modality introduced by the double-cover of $\mathcal{S}(3)$ (see section A.1.6). For delta actions, the local tangent space representation significantly outperforms others and performs better than global actions. Attaching the tangent space to the local frame always puts the cut-locus on the farthest side from the current orientation and prevents it from impacting uniqueness and continuity. All results are shown in figures 16 and 17.
@@ A.2.3 TD3 Results
A.3 Implementations

### v2

@@ Abstract
Many robotic control tasks require policies to act on orientations, yet the geometry of $\mathrm{SO}(3)$ makes this nontrivial. Because $\mathrm{SO}(3)$ admits no global, smooth, minimal parameterization, common representations such as Euler angles, quaternions, rotation matrices, and Lie algebra coordinates introduce distinct constraints and failure modes. While these trade-offs are well studied for supervised learning, their implications for actions in reinforcement learning remain unclear. We systematically evaluate $\mathrm{SO}(3)$ action representations across three standard continuous control algorithms, PPO, SAC, and TD3, under dense and sparse rewards. We compare how representations shape exploration, interact with entropy regularization, and affect training stability through empirical studies and analyze the implications of different projections for obtaining valid rotations from Euclidean network outputs. Across a suite of robotics benchmarks, we quantify the practical impact of these choices and distill simple, implementation-ready guidelines for selecting and using rotation actions. Our results highlight that representation-induced geometry strongly influences exploration and optimization and show that representing actions as tangent vectors in the local frame yields the most reliable results across algorithms. The project webpage and code are available at amacati.github.io/so3_primer.
@@ 1 Introduction
What is still missing is a general, systematic evaluation of $\mathrm{SO}(3)$ representations in deep reinforcement learning (RL). While the intuitions and results from prior studies on input/observation representations for orientations also apply to the RL setting, the most suitable $\mathrm{SO}(3)$ action representation remains unclear. Action representation requires special attention, as it shapes the exploration dynamics induced by stochastic policies and exploration noise, and has implications for action clipping to comply with actuation constraints. Prior works have proposed specific action representations for a narrow set of problems and algorithms (Alhousani et al., 2023a; b). Schuck et al. (2025a) recently attempted to tackle the general question of best action and observation representations for RL, but limited their investigation to Deep Deterministic Policy Gradients (DDPG) under sparse rewards.
@@ 2.1 The SO(3) Manifold
Representation $\Delta$ Action Dim. Cover Smooth Singularities Constraints
Matrix $\bm{R}$} $\Delta\bm{R}$ 9 single + - ${\bm{R}}^{T}{\bm{R}}=\bm{I}$ $\det{\bm{R}}=1$
Quaternion $\bm{q}$} $\Delta\bm{q}$ 4 double + - ${\bm{q}}||_{2}=1$
Euler angles $(\phi,\theta,\psi)$ $\Delta(\phi,\theta,\psi)$ 3 multi - +
Lie algebra $\mathfrak{m}$ ($^{\mathcal{E}}\bm{\tau}$) $^{\bm{s}}\bm{\tau}$ 3 multi - +
@@ 2.4 Projections
For TD3’s deterministic policies, this suffices to guarantee actions lie in $\mathrm{SO}(3)$. Stochastic policies, on the other hand, are often parameterized as multivariate Gaussians. Applying the projection to each sampled action guarantees on-manifold actions, but it also warps the action distribution and renders log probabilities intractable. PPO and SAC rely on accurate log probabilities, and closed-form corrections for normalization on $\mathcal{S}(3)$ or SVD-based projections are not readily available.
In this paper, we adopt a practical compromise: we project the mean inside the network wherever possible and sample in the ambient Euclidean space. The sampled, off-manifold actions are projected again within the environment to a valid rotation. This approach keeps training compatible with standard log-probability computations while ensuring feasibility at execution time.
2.4.1 Unit-rotation Centering
Policy neural networks for RL are designed to initially produce zero-centered output distributions or actions. Counterintuitively, the projections for quaternion and matrix representations map outputs around zero to a wide range of rotation actions (see section A.1.1). This is particularly relevant for delta actions, as the agent must first learn the unit rotation that prevents it from rotating.
One possible remedy is a custom policy network that adds the unit rotation to the action mean. We discuss the performance impact on delta matrix and quaternion representations in section 3.3, and provide ablations on corrections as well as more advanced benchmarks in section A.3. Local tangent vectors and delta Euler angles center around the unit operation and are thus unaffected.
@@ 3.1 Environment Setup
with the geodesic distance $d({\bm{R}}_{1},{\bm{R}}_{2})=\arccos\Big(\tfrac{\mathrm{tr}({\bm{R}}_{1}^{\top}{\bm{R}}_{2})-1}{2}\Big).$ Intuitively, equation 3 takes the shortest path towards the desired orientation ${\bm{R}}_{a}$ with a maximum step length of $\alpha_{max}$. Figure 1 visualizes the environment dynamics. Dense rewards $r_{t}^{\mathrm{dense}}=-d({\bm{R}}_{t},{\bm{R}}_{g})$ are defined as the negative angle to the goal. Sparse rewards are 0 when the angle between state and goal $d({\bm{R}}_{t},{\bm{R}}_{g})\leq 0.1$ and -1 everywhere else. Termination occurs after a fixed step limit of $50$.
@@ 3.2 Performance Comparison
We benchmark PPO, SAC, and TD3 in the pure-rotation environment using four action parameterizations: rotation matrices, unit quaternions, tangent-space (rotation vectors), and Euler angles, each evaluated as global and delta actions. All other components remain fixed: network architectures, training budgets, observation spaces, and reward definitions are identical across conditions. The results are presented in table 2, with the best results highlighted in blue, and the second-best shown in bold. Results are averaged over 50 runs each. See section A.2 for the training curves, section A.4 for a discussion on Zhou et al. (2019)’s representation, and section A.8 for hyperparameters.
@@ 3.3 Explaining the Effects of SO(3) Action Representations
Conclusion Uniqueness and smoothness benefit learning, but the properties do not have to hold globally for agents with limited angular step sizes. What matters is that discontinuities and singularities are out of reach for the action space, e.g., as in the local tangent space.
@@
Figure 2: Distribution of Euler angles sampled from $\mathcal{N}(0,2)$ and squashed with $\tanh$. The resulting orientations are visualized in the 3D Lie algebra $\mathfrak{m}$. Note that samples also lie within the spherical space.
@@
Conclusion The projection onto $\mathrm{SO}(3)$ warps common exploration distributions, which significantly impacts convergence. Euler angles and quaternions are most affected by this, matrices to some degree, and local tangent spaces least.
@@
Conclusion Increased entropy regularization drives actions to larger magnitudes, but fails to increase the diversity of matrix and quaternion actions. It influences exploration, particularly in sparse reward environments, but does not change the final policy. Attractions to Euler singularities only become relevant with extremely high entropy bonuses.
Hypothesis 4 Unit rotation-centered policies improve performance for delta actions.
Figure 3: PPO learning curves for unit rotation-centered delta actions using delta quaternions.
Section 2.4.1 remarks that due to the projection onto $\mathrm{SO}(3)$, small, zero-centered actions can result in a large spread of rotations for quaternion and matrix representations. Centering the output around the unit rotation should lead to less erratic initial exploration. In addition, agents do not have to learn the group-specific representation of the unit operation, i.e., unit quaternions and rotation matrices.
We test the hypothesis by adding the constant unit rotation to the policy mean. While Figure 3 shows a clear performance improvement over the non-centered delta actions, we find that this mostly holds for PPO, and the same modification produces mixed results for TD3 and SAC. For details on the implementation and the complete set of results, including benchmarks from section 4, refer to section A.3.
Conclusion Centering delta actions around the unit rotation improves the performance of quaternion and matrix representations in PPO. For SAC and TD3, the results are not conclusive. Delta tangent and Euler actions are unaffected as they are centered by construction.
Hypothesis 5 Scaling actions to the range of permissible rotations boosts performance. As outlined in section 2.5, tangent vector increments in the local frame can easily be scaled. Restricting the network output to the range of possible rotations should be more efficient because the agent does not need to learn that actions with the same direction and magnitudes larger than $\alpha_{max}$ lead to the same outcome. Furthermore, it removes discontinuities at the cut locus from the action space and allows more fine-grained control in the relevant action ranges.
We can see the effects in all three algorithms. Ablation studies with unscaled tangent increments and dense rewards consistently exhibit a performance difference of around 1.5-1.5 compared to scaled tangent vectors across PPO, SAC, and TD3. The effect is smallest for PPO, where the policy initialization around zero prevents frequent sampling of unscaled actions that reach the cut locus. Agents trained on sparse rewards exhibit a similar decrease in performance. More importantly, however, sparse reward agents sometimes take longer to converge to a near-optimal policy in TD3, and in rare cases show degraded performance in SAC (see section A.1.5). SAC’s failures are caused by the entropy bonus driving actions to regions of the representation with discontinuities and singularities as outlined above.
Conclusion Scaling tangent vectors to the range of permissible angles improves performance and stability. For PPO, pay special attention to the log standard initializations.
@@ 4 Robotic Benchmarks
In real-world problems, rotation actions in $\mathrm{SO}(3)$ rarely act on purely rotational dynamics as in section 3. Instead, they appear in a larger context, such as controlling a robot arm’s pose and gripper to grasp objects, or controlling a drone’s orientation and total thrust for trajectory tracking. In this section, we validate the relevance and transferability of our findings by applying the different action parameterizations on three robot benchmarks. Here, we focus on the most promising action combinations, i.e., global matrix and quaternion representations, and delta tangent and Euler angles.
@@
In figure 4, action representations significantly impact convergence speed on both benchmarks. Actions in the local tangent space consistently outperform other representations by converging faster and achieving higher rewards. Surprisingly, Euler angles are second before rotation matrices and quaternions. The reason is the limited range of angles required in the tasks. The drone cannot deviate too much from the upright orientation without crashing, and hence the policy remains in a region of $\mathrm{SO}(3)$ where Euler angles are still well-behaved. In contrast, absolute quaternion and matrix actions are highly random at initialization (see section A.1.1), which leads to fast crashes and limits progress. Ablations in section A.3 show that for these unstable environments, unit-centering (see section 2.4.1 and hypothesis 4 in section 3.3) is a key factor for performance. It is highly advisable to use relative and unit-centered representations this class of systems.
@@ 5 Conclusion
One limitation of this paper is its restriction to state-based observations and small networks. Observations do not affect the action space; thus, our results likely still apply. However, this requires empirical evidence. In addition, we did not consider discrete action space algorithms, which open up entirely new questions, e.g. on discretization schemes and the required density of the SO(3) cover. We also noticed a lack of suitable benchmarks that require control over the full $\mathrm{SO}(3)$ manifold. Our extension to the HER environments could act as a starting point to build up a standard set of tasks focusing on this capability. Finally, diffusion policies have found widespread adoption for imitation learning in robotics. A similar study on suitable representation choices for diffusion may reach significantly different conclusions due to diffusion models’ multi-modality and noise processes.
@@ 5.1 Reproducibility Statement
The code to reproduce our experiments is publicly available at github.com/amacati/so3_primer. It includes the training scripts and instructions for running the experiments we describe.
5.1.1 Acknowledgments
We thank SiQi Zhou and Marcel Rath for their helpful feedback on the manuscript, and Jan Brüdigam for the discussions on SO(3) representations and their impact on RL training. This work was supported by the Robotics Institute Germany under BMBF grant 16ME0997K, and by the Humboldt Professorship for Robotics and Artificial Intelligence.
@@ A.1.1 Projections of Noise Samples
For stochastic policies used in PPO and SAC, a natural question is whether projecting sampled actions to the valid representation manifold speeds up convergence, as the critic never sees off-manifold actions. Ablation runs in figure 8 show this is false. Projections lead to a significant performance loss in PPO, because the probability ratios $\tfrac{\pi_{\theta}(\mathbf{a}\mid\mathbf{s})}{\pi_{\theta_{old}}(\mathbf{a}\mid\mathbf{s})}$ used in its clipped surrogate objective no longer matches the ratio of the unprojected action. Newly computed probabilities use projected actions, while the stored probabilities of the old policy are based on non-projected actions. SAC remains unaffected as action probabilities are re-computed online during policy updates, with projections applied afterwards. However, we do not observe any performance improvements.
@@
Based on our ablations, we recommend not projecting the actions of PPO and SAC after sampling, and instead relying on the environment for the action projection. Introducing projections offers no performance gains in the best case, and has the potential to severely harm convergence.
@@ A.1.5 Scaling Tangent Vectors
Section 2.5 outlines how tangent vectors can be scaled to only encompass the range of possible rotations with a limited angle $\alpha_{max}$. In section 3.3, we explain how this helps avoid the cut locus of the tangent space at large action norms. Here, we present the training curves of our ablations in figures 10, 11 and 12. As previously stated, the final mean reward of scaled policies is slightly increased by $1.5$ to $2$ due to more fine-grained control. Much of the unscaled policies’ action space lies outside the $\alpha_{max}$ limit and thus maps to the same actions.
@@
The effect is smallest for PPO, where the initialization of the policy leads to actions close to zero. For SAC and TD3, the effect is slightly more visible in dense environments. Agents in sparse environments infrequently converge with significant delay (TD3) or fail to converge (SAC), causing the increase in variance across runs. The entropy bonus drives actions towards larger magnitudes around the cut locust before learning a reasonable policy, which causes the complete failures in SAC. Dense rewards prevent this by balancing the entropy bonuses with a continuous signal towards successful policies from the beginning.
@@ A.1.7 Extended Results for Entropy Regularization in PPO and SAC
A.2 Full Experimental Results on the Idealized Environment
@@ A.2.2 SAC Results
For dense rewards and global actions, the rotation matrix representation converges fastest. We attribute this to the uniqueness and smoothness of the representation. Quaternions are second with decreased convergence stability because of the multi-modality introduced by the double-cover of $\mathcal{S}(3)$ (see section A.1.6). For delta actions, the local tangent space representation significantly outperforms others and performs better than global actions. Attaching the tangent space to the local frame always puts the cut-locus on the farthest side from the current orientation and prevents it from impacting uniqueness and continuity. All results are shown in figures 17 and 18.
@@ A.2.3 TD3 Results
A.3 Extended Discussion of Unit Rotation-Centered Delta Actions
Section 2.4.1 discusses that policy networks using quaternion and matrix representations are not centered around the unit rotation for delta actions, potentially leading to worse performance. To address this, the network output can be modified such that a zero output corresponds to the identity rotation $\bm{I}$ (or unit quaternion $\bm{q}_{\mathbf{I}}$). We investigate two distinct implementations of this centering strategy for quaternion and matrix representations and evaluate their impact on learning stability and final performance.
A.3.1 Comparisons of Implementations
We denote the raw output of the neural network as $\bm{q}_{\text{net}}$ and $\bm{R}_{\text{net}}$, respectively. We define two variants for mapping this output to a pre-projection action $\bm{q}_{\text{pre}}$ and $\bm{R}_{\text{pre}}$, which are subsequently projected to obtain the valid rotation.
Additive Bias.
The simplest approach is to add the identity element to the network output.
$\bm{q}_{\text{pre}}$ $=\bm{q}_{\text{net}}+\bm{q_{I}}$
$\bm{R}_{\text{pre}}$ $=\bm{R}_{\text{net}}+\bm{I}$
This variant ensures that if $\bm{q}_{\text{net}}$ or $\bm{R}_{\text{net}}\approx 0$, the resulting action is close to the identity. However, this method limits the set of reachable actions. Assuming $\tanh$ activations bound $a_{\text{net}}$ within [1,1][-1,1], the scalar component of $\bm{q}_{\text{pre}}$ is restricted to the range $[0,2]$. Consequently, the policy cannot represent the conjugate unit quaternion $\bm{q_{I}}^{*}$ directly. A similar property holds for the matrix representation.
Scaled Bias.
To mitigate the range limitation of the additive variant, we can also implement the mapping with a scaling factor dependent on the identity element.
$\bm{q}_{\text{pre}}$ $=\bm{q}_{\text{net}}\odot(\bm{q_{I}}+\bm{1})+\bm{q_{I}}$
$\bm{R}_{\text{pre}}$ $=\bm{R}_{\text{net}}\odot(\bm{R_{I}}+\bm{1})+\bm{I}$
Here, $\odot$ denotes element-wise multiplication and $\bm{1}$ is a tensor of ones with matching dimensions. By scaling the network output, the policy can reach all possible quaternion configurations, including the negative hemisphere. However, this introduces non-uniform sensitivities, since the scalar component $w$ and the diagonal elements of the rotation matrix, respectively, induce larger changes in the resulting rotation magnitude compared to the additive variant.
A.3.2 Results on the Idealized Environment
We benchmark both implementations against the standard, uncentered parameterizations in our idealized environment. Results are averaged over 50 runs as in the main text.
Figure 21: PPO learning curves for unit rotation-centered delta actions using delta quaternion (left) and matrix (right) action representations in the idealized environment.
For PPO, the modified relative actions provide a clear benefit. Both converge slightly faster than their absolute variants and significantly improve over the uncentered delta actions (see figure 21). Comparing the two variants, the additive mapping achieves a slight performance advantage over the scaled one. For quaternions, only the additive variant achieves similar performance to the absolute representation. For the matrix representation, both variants converge to a lower performance than the absolute representation.
Figure 22: SAC learning curves for unit rotation-centered delta actions and dense rewards using delta quaternion (left) and matrix (right) action representations in the idealized environment.
The results for SAC present a contrast to PPO. As can be seen in figure 22, in dense rewards, the scaled variant slightly improves over regular delta and absolute actions for quaternions, whereas the additive one leads to significantly worse performance. The matrix representations are almost unchanged compared to the default implementation. In sparse rewards (see figure 23), the additive version performs slightly worse compared to the default delta orientation, whereas the scaled version slightly outperforms it. None of the three, however, yields a successful policy. The same is true for the rotation representation.
Figure 23: SAC learning curves for unit rotation-centered delta actions and sparse rewards using delta quaternion (left) and matrix (right) action representations in the idealized environment.
Figure 24: TD3 learning curves for unit rotation-centered delta actions and dense rewards using delta quaternion (left) and matrix (right) action representations in the idealized environment.
TD3 shows a similarly mixed picture in figure 24, where both variants slightly underperform with dense rewards compared to the default quaternion implementation, and have significantly increased variances between runs. The matrix representation shows no real difference between the default implementation and the improved versions. In sparse rewards, both modified quaternion representations outperform the baseline, but still exhibit significant variance and fall short of the consistency and performance of the absolute representation. The modified matrix representations in figure 25 significantly outperform the baseline and approach the performance of the absolute representation, but have a higher variance between runs.
Figure 25: TD3 learning curves for unit rotation-centered delta actions and sparse rewards using delta quaternion (left) and matrix (right) action representations in the idealized environment.
Our experiments indicate that while unit-centering addresses some of the issues associated with delta rotations, its practical impact varies by algorithm. The performance difference is negligible for many configurations. The notable exceptions are PPO, where centering is crucial for competitive performance of delta actions, and SAC, where the scaled variant is favorable for quaternions. In the specific case of PPO, we observe a slight advantage of the additive variant over the scaled one.
Revisiting Unstable Systems with Unit-biased Relative Actions.
In section 4, we hypothesized that zero-centered actions are particularly beneficial for unstable robotic systems, such as drones, where maintaining the initial attitude is safer than a random rotation. Centering actions around the unit rotation should theoretically allow quaternion and matrix representations to stabilize and converge faster on these tasks.
Since the additive variant performed best for PPO, we selected it to validate this hypothesis on the trajectory tracking and drone racing benchmarks. We compare this centered variant against the global action baseline for both quaternion and matrix actions. For a fair comparison, we retune the hyperparameters using the same sweep settings as for the baseline.
Figure 26: PPO learning curves for the drone trajectory tracking environment using delta quaternion (left) and matrix (right) action representations with unit rotation bias.
Our results confirm the hypothesis. The modified relative quaternion actions in figure 26 demonstrate dramatically improved convergence speed compared to uncentered variants. Similar holds true for matrices, although neither manage to outperform the local tangent representation. By initializing the policy output effectively at the identity rotation, the drone is less prone to erratic initial behavior that leads to immediate crashes, thereby allowing the agent to gather meaningful experience earlier in the training process.
Figure 27: PPO learning curves for the drone racing environment using delta quaternion (left) and matrix (right) action representations with unit rotation bias.
While less dramatic, our results on the drone racing task in figure 27 clearly show the same trend. We conclude that for unstable systems, the unit-centering of representations is a key influence on performance.
A.4 Comparison of Rotation Matrix and 6D Representations
Beyond the standard parameterizations discussed in the main text, several previous studies have explored suitable rotation representations for learning. One influential representation, particularly noted for its properties in optimization and supervised learning contexts, is the continuous 6D representation proposed by Zhou et al. (2019).
The idea behind this representation is that a full 9-element rotation matrix is over-parameterized. Defining two orthogonal column vectors is sufficient to reconstruct the complete orientation without losing the uniqueness and smoothness of the representation. Since a rotation matrix belonging to $\mathrm{SO}(3)$ describes a right-handed coordinate system, the third axis is entirely defined by the cross product of the first two. Implementation-wise, the policy outputs a 6D vector $x\in\mathbb{R}^{6}$ that is treated as two raw 3D vectors. To ensure orthogonality and unit norm, we then apply the Gram-Schmidt process to these vectors.
While this representation enjoys a lower dimensionality than the standard flattened rotation matrix, Geist et al. (2024) point out that it empirically falls short of the standard SVD-projected matrix representation on several tasks. Since the arguments regarding continuity and smoothness apply similarly to both the 6D representation and standard matrices, and given that rotation matrices are more prevalent in standard robotics pipelines, we selected matrices for our primary comparison. In this section, we provide a direct comparison to demonstrate that the performance of the 6D representation is indeed comparable to, but rarely strictly better than, the standard matrix representation in reinforcement learning settings.
A.4.1 Results on the Idealized Environment
We benchmark the 6D representation against the standard matrix representation across all algorithms with 50 runs per variant as in the main text.
Figure 28: PPO learning comparison between the 6D representation and rotation matrices in the idealized environment for both absolute (straight lines) and delta (dotted lines) actions.
For PPO, the 6D representation slightly outperforms the standard matrix representation in terms of both convergence speed and final performance when using absolute actions. Conversely, for delta representations, the 6D parameterization falls short of the matrix baseline (see figure 28).
Figure 29: SAC learning comparison between the 6D representation and rotation matrices in the idealized environment for absolute (straight lines) and delta (dotted lines) actions on dense (left) and sparse (right) rewards.
The results for SAC, visualized in figure 29, present a similar picture. In the dense reward setting, the performance of relative and absolute representations is practically indistinguishable between the two parameterizations. In the sparse reward setting, we observe a minor improvement for the 6D representation across both absolute and relative modes. However, it is important to note that neither the matrix nor the 6D representation successfully learns a high-quality policy under these sparse reward conditions, sharing the same failure modes and training characteristics.
Figure 30: TD3 learning comparison between the 6D representation and rotation matrices in the idealized environment for absolute (straight lines) and delta (dotted lines) actions on dense (left) and sparse (right) rewards.
In TD3, the standard matrix variants slightly outperform their 6D counterparts generally. The only meaningful difference exists between the delta matrix and the delta 6D actions under sparse rewards, where the matrix representation converges to a better policy on average (see Figure 30). However, consistent with our main results, both representations yield relatively poor policies in this setting.
Overall, the performance curves of matrix and 6D representations are very similar across runs. PPO with absolute actions shows a slight preference for the 6D representation, whereas TD3 generally favors the standard matrix representation. A possible explanation for this result is the increased robustness of the full matrix representation to large disturbances such as the uniform exploration noise used in TD3, compared to the Gram-Schmidt reconstruction used in the 6D mapping. However, we emphasize that these performance differences are relatively minor and should not be overinterpreted as a definitive advantage for either representation.
Robot Benchmarks for PPO
Since the 6D representation slightly outperformed its matrix counterpart on PPO (see figure 28), we rerun the robot benchmarks that use PPO for trajectory following and drone racing.
Figure 31: PPO learning comparison between the 6D representation and rotation matrices on the trajectory following task (left) and on drone racing (right).
The results in figure 31 display an inconclusive picture. While the 6D representation improves notably on the trajectory following task, the deviations between runs are large. Other effects such as unit-centering of delta actions (see section A.3) are significantly more consistent in improving convergence. The results on racing are again in line with our intuition that the two representations perform very similar. Although matrix marginally outperforms the 6D representation, the difference is not significant.
A.5 Implementations